OpenStudy (anonymous):

does anyone know how to solve: x+y=8, xy=15

OpenStudy (anonymous):

8yx-15xy

OpenStudy (anonymous):

change x+y=8 to x=8-y and then subtitute to xy=15, (8-y)y=15, 8y-y^2=15, y^2-8+15=0, y=-3|-5. when y=-3 so x-3=8, x=11, x+y=8 => 11-3=8 but xy=15=>11.-3 !=15, when y=-5, x-5=8 => x=13 but xy=15, 13.-5 !=15. So the best answer is x=3 and y=5 or x=5 and y=3